OpenStudy (anonymous):

Can Someone help me? Wlil give Medal. Please no direct answers. 8 + 9y x 4y x 3 = ?

OpenStudy (anonymous):

@Atsie @mathmale @goopyfish908 Plz help

pooja195 (pooja195):

\[\huge~\rm~8 + 9y \times 4y \times 3 = ?\] like this?

pooja195 (pooja195):

or are they x's?

OpenStudy (anonymous):

yeah

OpenStudy (mathmale):

Except in very elementary math, it's best NOT to use " x " to denote multiplication. Instead, use " * " or enclose the multiplicands within parentheses: (9x)(8y) = 72 xy

OpenStudy (anonymous):

@pooja195 They are multiplied

OpenStudy (anonymous):

Sorry didnt think you would understand @mathmale But thanks :)

OpenStudy (anonymous):

@mathmale Here you go, a medal :D

pooja195 (pooja195):

Ok let's start by multiplying the y's \[\huge~\rm~ 9y~\times 4y=?\]

OpenStudy (anonymous):

That is 72y

pooja195 (pooja195):

How did you get that?

OpenStudy (anonymous):

9 * 7 = 72 then just add y " 72y "

OpenStudy (texaschic101):

whats 9 * 4 and whats y * y ?

pooja195 (pooja195):

@Gabe325 where did you get the 7 from ?

OpenStudy (anonymous):

CUZ 7 * 9 = 72

pooja195 (pooja195):

Where did you get 7 from i don't see it anywhere in the equation you gave

OpenStudy (anonymous):

nevermind..

pooja195 (pooja195):

Ok multiply \[\huge~\rm~ 9y~\times 4y=?\] (9)(4) =?

OpenStudy (anonymous):

72!

pooja195 (pooja195):

9 * 4 does not equal 72.Try again. Go over your 9 times table. 9 x 1= 9 9 x 2=18 9 x 3=27 9 x 4=?

OpenStudy (anonymous):

oops, its 32

OpenStudy (anonymous):

36*

pooja195 (pooja195):

Good y x y=?

OpenStudy (anonymous):

dunno.

pooja195 (pooja195):

You have 2 y's it would be y^2 now we have \[\huge~\rm~8 + 36 y^2\times3\]

pooja195 (pooja195):

\[\huge~\rm~36 \times 3=? \]

OpenStudy (anonymous):

108

pooja195 (pooja195):

good you get 108y^2 and you're left with positive 8 you can't combine it because it is not a like term.

OpenStudy (anonymous):

yep

pooja195 (pooja195):

Yeah that's all it is :)

OpenStudy (anonymous):

so answer is 108^2?

pooja195 (pooja195):

108y^2+8

OpenStudy (anonymous):

that is 116^2?

pooja195 (pooja195):

You don't combine it they are not like terms

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

oh ok, that is one of the answers

OpenStudy (anonymous):

108^2 + 8

OpenStudy (anonymous):

yay! i gots 100%

OpenStudy (anonymous):

To reward you, i give medal ;D

pooja195 (pooja195):

100% on what? .-.

OpenStudy (anonymous):

Paper

pooja195 (pooja195):

oh ok =) well yw ^_^

OpenStudy (anonymous):

Thanks for medal :)

pooja195 (pooja195):

=)

OpenStudy (anonymous):

Thanks for medal :)

OpenStudy (anonymous):

Ok bye :D